OpenStudy (anonymous):

9 squared * 9 negative squared without negative exponents.

OpenStudy (mertsj):

\[9^2(9)^{-2}=\frac{9^2}{9^2}=1\]

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw